# bibliography.bib

### v2

@@ 2.2.2 Positivity of the Lower Lipschitz Constant
    $00<\lVert a-\delta\rVert=\lim_{n\to\infty}\lVert s_{n}a_{n}-s_{n}e_{n}\rVert\leq\lim_{n\to\infty}\lVert s_{n}a_{n}-s_{n}U_{g_{1}}e_{n}\rVert=$ 0 .
@@
Starting from vector $z_{1}$ and the sequences $(x_{n})_{n}$, $(y_{n})_{n}$ observed in 2.1 after $d-1$ steps of algorithmic construction of part (2) of 2.12 we get $d$ non-zero vectors $\{z_{1},\dots,z_{d}\}$ and a pair of sequences $(\tilde{x}_{n})_{n}$, $(\tilde{y}_{n})_{n}$ such that
@@
Recall that from part (2) of 2.6 we have that $\exists M_{0}\in\mathbb{N}$ such that $\forall n\geq M_{0}$ and $(i,j)\in S$
@@
00 0 $=\lim_{n\to\infty}\frac{\lVert\Phi_{{\textbf{w}},S}(\tilde{x}_{n})-\Phi_{{\textbf{w}},S}(\tilde{y}_{n})\rVert^{2}}{\operatorname{{\mathbf{d}}}([x_{n}],[y_{n}])^{2}}$
@@
$\Phi_{{\textbf{w}},S}(\sum_{r=1}^{d}z_{r})=\Phi_{{\textbf{w}},S}(\sum_{r=1}^{d}z_{r}+\epsilon\tilde{z})$
which contradict the injectivity property. 1.4 is now proved.
@@ 3 Dimension reduction using linear maps
The idea of using dimension reduction linear maps goes back many years. The famous Johnson-Lindenstrauss Lemma [johnson1986extensions] provides a nearly isometric projection for finite metric spaces. More recently, [Cahill19] shows that a dimension reduction linear map preserves Lipschitz properties for translation invariant polynomial embeddings.
The first result that combines sorted co-orbits with linear maps was shown in [balan2022permutation] in the case of the group of permutations $G=S_{n}$ acting by left multiplication on $n\times d$ real matrices, ${\mathscr{V}}=\mathbb{R}^{n\times d}$. In that particular case, the dimension of the intermediary co-orbit space was $n(1+(d-1)n!)$ while the final target space was shown to be $2nd=2\,dim({\mathscr{V}})$. The authors of [dym2022low] show that intermediate embedding can be realized in dimension $(2nd+1)n$ instead of $n(1+(d-1)n!)$ with the final target space of dimension $12nd+$ 1. A more careful analysis of such an embedding by Matthias Wellershoff [Matthias2023a] proved that $12nd+$ 1 can be replaced by $2nd-d$. In the Ph.D. thesis [tsoukanis2024g], one of the authors showed that the decrease of dimension from $2nd$ to $2nd-d$ represents the dimension of a certain joint eigenspace of this representation.
The exact statement of our dimension reduction result is included in 1.6.
To prove 1.6 we need first to construct a few objects. Let $D:{\mathscr{V}}\times{\mathscr{V}}\to\mathbb{R}^{m}$ be the nonlinear map $D(x,y)=\Phi_{{\textbf{w}},S}(x)-\Phi_{{\textbf{w}},S}(y)$. Its range $E$ is defined by $E=\operatorname{Ran}(D):=\{\Phi_{{\textbf{w}},S}(x)-\Phi_{{\textbf{w}},S}(y):x,y\in{\mathscr{V}}\}=\operatorname{Ran}(\Phi_{{\textbf{w}},S})-\operatorname{Ran}(\Phi_{{\textbf{w}},S})$.
@@
Let $W_{a}=\operatorname{span}\{w_{1,a},\dots,w_{m-r,a}\}$. Notice that each $\operatorname{span}(w_{1,a},\dots,w_{m-r,a}\}$ is a linear subspace of $\mathbb{R}^{m}$ of dimension $m-r$. If $r\geq 1$, each $W_{a}$ is a proper subspace of $\mathbb{R}^{m}$.
A generic $v_{1}\in\mathbb{R}^{m}$ satisfies $v_{1}\not=0$ and $v_{1}\notin\cup_{a=1}^{N}W_{a}$. Replace each $W_{a}$ with $W_{a}^{1}=\operatorname{span}(W_{a},\{v_{1}\})$, subspaces of dimension $\dim(W_{a}^{1})=\dim(W_{a})+1=m-r+1$. If $\dim(W_{a}^{1})<m$, repeat this process inductively and obtain $v_{2},\dots,v_{r}$ until $\dim(W_{a}^{r})=m$. The procedure produces a set of vectors $(v_{1},\dots,v_{r})$ that satisfy the condition $\prod_{a}R_{a}(v_{1},\dots,v_{r})\not=0$. Hence $\mathbb{U}\neq\emptyset$. This ends the proof of 3.1. ∎
@@
If $m\leq 2d-d_{G}$ then the conclusion is satisfied for any full-rank $\ell$. Therefore assume $m>2d-d_{G}$. A generic linear map $\ell:\mathbb{R}^{m}\to\mathbb{R}^{2d-D_{G}}$ is full-rank. Hence $\dim(\operatorname{Ran}(\ell))=2d-d_{G}$, and thus $\dim(\ker(\ell))=m-2d+d_{G}$. On the other hand, for a generic linear map $\ell$ 3.1 with $r=m-2d+d_{G}$, $s=2d-d_{G}$, and $T=(N!)^{2p}$, implies
@@
3.2 shows that, a generic linear map $\ell:\mathbb{R}^{m}\rightarrow\mathbb{R}^{2d-d_{G}}$ satisfies $\ker(\ell)\cap\operatorname{Ran}(D)=\{0\}$. Thus, if $x,y\in{\mathscr{V}}$ so that $\Psi_{{\textbf{w}},S,\ell}(x)=\Psi_{{\textbf{w}},S,\ell}(y)$ then $\ell(D(x,y))=0$. Therefore $D(x,y)=0$. Since $\hat{\Phi}_{{\textbf{w}},S}$ is injective it follows $x\sim y$. Thus, $\hat{\Psi}_{{\textbf{w}},S,\ell}$ is also injective.
From 1.4 we have that, if the map $\hat{\Phi}_{{\textbf{w}},S}$ is injective then it is also bi-Lipschitz. Let $a\leq b$ denote its bi-Lipschitz constants.
@@
Finally from 3.2 and 3.3 with $r=m-2d+d_{G}$, $s=2d-d_{G}$, $T=(N!)^{2p}$, $F_{a}=Ran(L_{\pi_{1},\dots,\pi_{2p}})$ we have that for a generic linear map $\ell$, for all $x,y\in{\mathscr{V}}$,
@@ 4 Universality of representation
In previous sections we constructed the embedding $\hat{\Phi}_{{\textbf{w}},S}$ of the quotient space $\hat{{\mathscr{V}}}$ into an Euclidean space $\mathbb{R}^{m}$. In this section we prove 1.7 and 1.8 that show that when $\hat{\Phi}_{{\textbf{w}},S}$ is injective, every continous or Lipschitz map $f:\hat{{\mathscr{V}}}\to H$ factors through $\hat{\Phi}_{{\textbf{w}},S}$.
The proof of 1.7 is based on Kirszbraun extension theorem [kirszbraun1934zusammenziehende] which is re-stated here for the reader’s convenience:
Theorem 4.1 (Kirszbraun extension theorem[kirszbraun1934zusammenziehende]).
@@
The second universality result 1.8 applies to the class of continuous functions (instead of Lipschitz functions).
The proof of 1.8 follows from the following extension of Tietze’s theorem [urysohn1925machtigkeit, dugundji1951extension].
Theorem 4.2 (Dugundji-Tietze [dugundji1951extension]).
@@
Proof of 1.8.
Let $S=\Phi_{{\textbf{w}},S}({\mathscr{V}})\subset\mathbb{R}^{m}$. Note that $S$ is a closed set as $\hat{\Phi}_{{\textbf{w}},S}$ is bi-Lipschitz and $\hat{{\mathscr{V}}}$ is complete. Let $t:S\mapsto L$ be defined by $t(\Phi_{{\textbf{w}},S}(x))=F(x)$, for all $x\in{\mathscr{V}}$. Note $t$ is continuous since $\hat{\Phi}_{{\textbf{w}},S}$ is bi-Lipschitz on $\hat{{\mathscr{V}}}$. By 4.2 there exists a continuous extension $T:\mathbb{R}^{m}\rightarrow L$ of $t$ that satisfies the convex hull property.
[…]

### v4

@@ 2.2.2 Positivity of the Lower Lipschitz Constant
    $0<\lVert a-\delta\rVert=\lim_{n\to\infty}\lVert s_{n}a_{n}-s_{n}e_{n}\rVert\leq\lim_{n\to\infty}\lVert s_{n}a_{n}-s_{n}U_{g_{1}}e_{n}\rVert=0.$
@@
Starting from vector $z_{1}$ and the sequences $(x_{n})_{n}$, $(y_{n})_{n}$ observed in ˜2.1 after d1d-1 steps of algorithmic construction of part (2) of ˜2.12 we get $d$ non-zero vectors $\{z_{1},\dots,z_{d}\}$ and a pair of sequences $(\tilde{x}_{n})_{n}$, $(\tilde{y}_{n})_{n}$ such that
@@
Recall that from part (2) of ˜2.6 we have that $\exists M_{0}\in\mathbb{N}$ such that $\forall n\geq M_{0}$ and $(i,j)\in S$
@@
$0$ $=\lim_{n\to\infty}\frac{\lVert\Phi_{{\textbf{w}},S}(\tilde{x}_{n})-\Phi_{{\textbf{w}},S}(\tilde{y}_{n})\rVert^{2}}{\operatorname{{\mathbf{d}}}([x_{n}],[y_{n}])^{2}}$
@@
[…]
which contradict the injectivity property. ˜1.4 is now proved.
@@ 3 Dimension reduction using linear maps
The idea of using dimension reduction linear maps goes back many years. The famous Johnson-Lindenstrauss Lemma [26] provides a nearly isometric projection for finite metric spaces. More recently, [25] shows that a dimension reduction linear map preserves Lipschitz properties for translation invariant polynomial embeddings.
The first result that combines sorted co-orbits with linear maps was shown in [10] in the case of the group of permutations $G=S_{n}$ acting by left multiplication on $n\times d$ real matrices, ${\mathscr{V}}=\mathbb{R}^{n\times d}$. In that particular case, the dimension of the intermediary co-orbit space was n(1+(d1)n!)n(1+(d-1)n!) while the final target space was shown to be $2nd=2\,dim({\mathscr{V}})$. The authors of [22] show that intermediate embedding can be realized in dimension $(2nd+1)n$ instead of n(1+(d1)n!)n(1+(d-1)n!) with the final target space of dimension $2nd+1$. A more careful analysis of such an embedding by Matthias Wellershoff [38] proved that $2nd+1$ can be replaced by 2ndd2nd-d. In the Ph.D. thesis [34], one of the authors showed that the decrease of dimension from $2nd$ to 2ndd2nd-d represents the dimension of a certain joint eigenspace of this representation.
The exact statement of our dimension reduction result is included in ˜1.6.
To prove ˜1.6 we need first to construct a few objects. Let $D:{\mathscr{V}}\times{\mathscr{V}}\to\mathbb{R}^{m}$ be the nonlinear map $D(x,y)=\Phi_{{\textbf{w}},S}(x)-\Phi_{{\textbf{w}},S}(y)$. Its range $E$ is defined by $E=\operatorname{Ran}(D):=\{\Phi_{{\textbf{w}},S}(x)-\Phi_{{\textbf{w}},S}(y):x,y\in{\mathscr{V}}\}=\operatorname{Ran}(\Phi_{{\textbf{w}},S})-\operatorname{Ran}(\Phi_{{\textbf{w}},S})$.
@@
Let $W_{a}=\operatorname{span}\{w_{1,a},\dots,w_{m-r,a}\}$. Notice that each $\operatorname{span}(w_{1,a},\dots,w_{m-r,a}\}$ is a linear subspace of $\mathbb{R}^{m}$ of dimension mrm-r. If $r\geq 1$, each $W_{a}$ is a proper subspace of $\mathbb{R}^{m}$.
A generic $v_{1}\in\mathbb{R}^{m}$ satisfies $v_{1}\not=0$ and $v_{1}\notin\cup_{a=1}^{N}W_{a}$. Replace each $W_{a}$ with $W_{a}^{1}=\operatorname{span}(W_{a},\{v_{1}\})$, subspaces of dimension $\dim(W_{a}^{1})=\dim(W_{a})+1=m-r+1$. If $\dim(W_{a}^{1})<m$, repeat this process inductively and obtain $v_{2},\dots,v_{r}$ until $\dim(W_{a}^{r})=m$. The procedure produces a set of vectors $(v_{1},\dots,v_{r})$ that satisfy the condition $\prod_{a}R_{a}(v_{1},\dots,v_{r})\not=0$. Hence $\mathbb{U}\neq\emptyset$. This ends the proof of ˜3.1. ∎
@@
If $m\leq 2d-d_{G}$ then the conclusion is satisfied for any full-rank $\ell$. Therefore assume $m>2d-d_{G}$. A generic linear map $\ell:\mathbb{R}^{m}\to\mathbb{R}^{2d-D_{G}}$ is full-rank. Hence $\dim(\operatorname{Ran}(\ell))=2d-d_{G}$, and thus $\dim(\ker(\ell))=m-2d+d_{G}$. On the other hand, for a generic linear map $\ell$ ˜3.1 with $r=m-2d+d_{G}$, $s=2d-d_{G}$, and $T=(N!)^{2p}$, implies
@@
˜3.2 shows that, a generic linear map $\ell:\mathbb{R}^{m}\rightarrow\mathbb{R}^{2d-d_{G}}$ satisfies $\ker(\ell)\cap\operatorname{Ran}(D)=\{0\}$. Thus, if $x,y\in{\mathscr{V}}$ so that $\Psi_{{\textbf{w}},S,\ell}(x)=\Psi_{{\textbf{w}},S,\ell}(y)$ then $\ell(D(x,y))=0$. Therefore $D(x,y)=0$. Since $\hat{\Phi}_{{\textbf{w}},S}$ is injective it follows $x\sim y$. Thus, $\hat{\Psi}_{{\textbf{w}},S,\ell}$ is also injective.
From ˜1.4 we have that, if the map $\hat{\Phi}_{{\textbf{w}},S}$ is injective then it is also bi-Lipschitz. Let $a\leq b$ denote its bi-Lipschitz constants.
@@
Finally from ˜3.2 and ˜3.3 with $r=m-2d+d_{G}$, $s=2d-d_{G}$, $T=(N!)^{2p}$, $F_{a}=Ran(L_{\pi_{1},\dots,\pi_{2p}})$ we have that for a generic linear map $\ell$, for all $x,y\in{\mathscr{V}}$,
@@ 4 Universality of representation
In previous sections we constructed the embedding $\hat{\Phi}_{{\textbf{w}},S}$ of the quotient space $\hat{{\mathscr{V}}}$ into an Euclidean space $\mathbb{R}^{m}$. In this section we prove ˜1.7 and ˜1.8 that show that when $\hat{\Phi}_{{\textbf{w}},S}$ is injective, every continous or Lipschitz map $f:\hat{{\mathscr{V}}}\to H$ factors through $\hat{\Phi}_{{\textbf{w}},S}$.
The proof of ˜1.7 is based on Kirszbraun extension theorem [27] which is re-stated here for the reader’s convenience:
Theorem 4.1 (Kirszbraun extension theorem[27]).
@@
The second universality result ˜1.8 applies to the class of continuous functions (instead of Lipschitz functions).
The proof of ˜1.8 follows from the following extension of Tietze’s theorem [35, 21].
Theorem 4.2 (Dugundji-Tietze [21]).
@@
Proof of ˜1.8.
Let $S=\Phi_{{\textbf{w}},S}({\mathscr{V}})\subset\mathbb{R}^{m}$. Note that $S$ is a closed set as $\hat{\Phi}_{{\textbf{w}},S}$ is bi-Lipschitz and $\hat{{\mathscr{V}}}$ is complete. Let $t:S\mapsto L$ be defined by $t(\Phi_{{\textbf{w}},S}(x))=F(x)$, for all $x\in{\mathscr{V}}$. Note $t$ is continuous since $\hat{\Phi}_{{\textbf{w}},S}$ is bi-Lipschitz on $\hat{{\mathscr{V}}}$. By ˜4.2 there exists a continuous extension $T:\mathbb{R}^{m}\rightarrow L$ of $t$ that satisfies the convex hull property.